# SLRQA: A Sparse Low-Rank Quaternion Model for Color Image Processing with Convergence Analysis

### v1

@@ 6.3.2 Color image denoising with NSS
(i) SLRQA-1
(j) SLRQA-2
(k) SLRQA-3
Figure 6: Color image denoising result of tested methods on “Image5” with $\tau=30$.
6.4 Color Image Inpainting
For color image inpainting problems, the compared methods include Bilinear Factorization (BF) methods such as LMafit [31], MCNMF [33], and Rank Minimization (RM) methods such as LADM [21], LRQA [10], and WNNM [13]555 The implementations that we use are available from:
LMafit: https://github.com/optsuite/LMaFit
MCNMF: https://xu-yangyang.github.io/codes/MCNF.zip
WNNM: https://www4.comp.polyu.edu.hk/cslzhang/code/WNNM_MC_code.zip
. We note that the BF methods require a rank estimate as a prior. For fairness, the LADM is extended to quaternion representation, which is referred to as QLADM and performs better than the LADM method. To fully demonstrate the performance of LMaFit and MCNMF, we choose the best PSNR and SSIM from the initial rank range [50, 80, 100, 120, 150].5080100120150[50,\,80,\,100,\,120,\,150].[ 50 , 80 , 100 , 120 , 150 ] .
The results of all tested algorithms are shown in Table 6. It is shown therein that SLRQA-NF-3 always enjoys the best performance compared with other methods in terms of PSNR and SSIM. The visual comparisons between SLRQA-NF and all competing inpainting methods on the Image2 with $\chi=0.7$ are shown in Figure 7 as a typical example. Although the LMaFit and MCNMF methods complete the image, the details of the recovered image are lost, and the recovered images are blurry. The images recovered by LRQA-1 and LRQA-2 methods have a relatively good effect, but the images are still slightly noisy or blurry. In addition, the images recovered by WNNM are noisy. In contrast, SLRQA-NF methods all demonstrate considerable inpainting results and have the highest PSNR and SSIM values.
(a) Image2
(b) Corrupted image
(c) WNNM
(d)LMaFit
(e) MCNMF
(f) QLADM
(g) LRQA-1
(h) LRQA-2
(i) SLRQA-NF-1
(j) SLRQA-NF-2
(k) SLRQA-NF-3
@@ 6.4 Color Image Inpainting
Image Algorithm LMaFit[31] WNNM [13] MCNMF [33] QLADM[21] LRQA-1 [10] LRQA-2 [10] SLRQA-NF-1 SLRQA-NF-2 SLRQA-NF-3
$\chi=0.5$ Image1 24.69/0.74 25.25/0.60 29.35/0.83 31.16/0.95 30.87/0.89 30.25/0.87 32.32/0.94 31.26/0.94 32.47/0.94
Image2 22.14/0.80 23.55/0.81 24.22/0.86 30.89/0.85 26.63/0.91 26.20/0.90 28.04/0.94 27.54/0.94 28.16/0.94
Image3 23.43/0.80 25.23/0.75 27.05/0.88 29.65/0.97 28.37/0.90 27.94/0.89 30.40/0.96 29.65/0.97 30.60/0.97
Image4 29.09/0.91 31.16/0.85 34.73/0.96 36.39/0.99 35.99/0.97 35.51/0.97 37.69/0.99 36.27/0.98 37.90/0.99
Image5 26.79/0.85 30.09/0.87 30.02/0.92 33.73/0.98 33.27/0.96 32.72/0.95 34.82/0.98 33.75/0.98 34.98/0.98
Image6 19.64/0.61 22.01/0.61 22.04/0.87 26.56/0.92 25.27/0.79 25.72/0.80 28.25/0.93 26.62/0.91 27.97/0.92
Image7 25.00/0.81 24.96/0.75 25.14/0.82 28.84/0.94 27.59/0.87 27.86/0.88 27.82/0.93 28.92/0.94 29.33/0.94
Image8 24.18/0.87 25.17/0.86 27.21/0.92 30.03/0.97 29.62/0.95 29.06/0.94 31.06/0.97 30.16/0.97 31.22/0.97
Image9 24.93/0.94 26.02/0.94 27.04/0.95 30.99/0.98 30.12/0.98 29.56/0.97 31.58/0.98 31.09/0.98 31.71/0.98
Image10 24.36/0.93 26.18/0.94 23.73/0.92 30.69/0.98 29.86/0.98 29.20/0.97 31.28/0.98 30.84/0.98 31.43/0.98
$\chi=0.7$ Image1 22.32/0.66 23.05/0.52 24.40/0.67 27.29/0.89 25.97/0.74 25.47/0.72 28.40/0.88 27.30/0.88 28.62/0.89
Image2 20.19/0.73 20.13/0.67 21.52/0.77 24.19/0.89 22.51/0.79 22.44/0.79 25.05/0.90 24.30/0.89 25.21/0.90
Image3 21.20/0.72 21.50/0.63 22.78/0.75 26.09/0.93 23.69/0.77 23.67/0.77 26.87/0.93 26.00/0.92 27.10/0.94
Image4 25.87/0.85 28.34/0.80 28.50/0.89 31.89/0.97 31.07/0.93 30.44/0.92 33.37/0.97 31.80/0.97 33.63/0.98
Image5 24.25/0.78 25.75/0.77 26.19/0.83 29.58/0.95 28.53/0.88 27.85/0.86 30.95/0.96 29.68/0.95 31.14/0.96
Image6 17.49/0.52 18.13/0.44 19.66/0.67 22.78/0.82 20.28/0.57 19.86/0.55 23.78/0.81 22.70/0.80 24.00/0.83
Image7 22.59/0.74 22.40/0.67 21.17/0.82 26.24/0.90 24.62/0.78 24.58/0.78 26.63/0.88 26.19/0.89 26.57/0.88
Image8 21.79/0.81 22.34/0.79 23.17/0.84 26.34/0.94 24.87/0.87 24.54/0.87 27.41/0.94 26.42/0.94 27.60/0.95
Image9 22.66/0.91 23.33/0.90 24.77/0.93 27.83/0.97 26.13/0.94 25.69/0.94 28.56/0.97 27.81/0.97 28.74/0.97
Image10 21.55/0.89 23.22/0.90 24.18/0.92 27.54/0.96 25.95/0.95 25.34/0.94 28.46/0.97 27.58/0.96 28.64/0.97
$\chi=0.8$ Image1 21.13/0.59 20.77/0.43 19.95/0.44 25.05/0.85 22.87/0.61 22.76/0.60 26.08/0.84 24.54/0.79 26.30/0.85
Image2 19.32/0.68 17.72/0.54 17.87/0.58 22.34/0.84 19.44/0.65 20.27/0.69 23.36/0.86 22.25/0.84 23.53/0.86
Image3 20.26/0.65 18.71/0.53 18.93/0.59 23.98/0.90 20.52/0.65 21.35/0.69 24.91/0.90 23.61/0.88 25.13/0.91
Image4 24.62/0.79 25.48/0.72 23.53/0.75 29.09/0.95 27.75/0.87 27.20/0.85 30.74/0.96 28.73/0.94 31.04/0.96
Image5 23.37/0.74 22.52/0.65 21.25/0.66 27.10/0.93 25.53/0.80 25.05/0.78 28.70/0.93 27.09/0.92 28.97/0.94
Image6 16.86/0.46 15.41/0.30 15.94/0.34 20.82/0.74 17.93/0.45 14.65/0.31 21.61/0.74 20.10/0.70 21.82/0.76
Image7 21.53/0.69 20.67/0.60 19.35/0.57 24.79/0.87 22.81/0.71 22.64/0.71 25.59/0.85 24.52/0.86 25.43/0.86
Image8 20.53/0.77 19.88/0.71 19.22/0.71 24.26/0.91 21.69/0.79 21.93/0.80 25.30/0.92 23.90/0.90 25.50/0.92
Image9 21.93/0.90 20.78/0.84 21.23/0.85 25.89/0.95 23.43/0.91 23.24/0.91 26.74/0.96 25.56/0.95 26.98/0.96
Image10 20.47/0.87 20.48/0.84 20.34/0.85 25.52/0.95 23.10/0.90 22.85/0.90 26.67/0.96 25.32/0.95 26.90/0.96
@@ 7 Conclusion
In this paper, we propose a novel SLRQA model for the color image process problems. Different from most existing models which only consider one or two properties from low-rankness, sparsity, and quaternion representation, SLRQA utilizes these properties all. Furthermore, SLRQA does not need an initial rank esamite. A PL-ADMM algorithm is proposed to solve the SLRQA model. Furthermore, we prove that any limit point of PL-ADMM for SLRQA is a stationary point under mild assumptions. In addition, the global convergence analysis of the sequence generated by PL-ADMM under the KL-inequality assumption is presented. When the observation is noise-free, a SLRQA-NF model of the limiting case of SLRQA model is proposed. Subsequently, a nonconvex PL-ADMM-NF algorithm is also proposed to solve the SLRQA-NF model. Extensive experiments for color image denoising and inpainting problems have demonstrated the robustness and effectiveness of SLRQA and SLRQA-NF.

### v2

@@ 5.4 Color Image Inpainting
(i) SLRQA-NF1
(j) SLRQA-NF2
(k) SLRQA-NF3
@@
To give a more detailed analysis, the errors of PL-ADMM-NF are also demonstrated for tested images. Figure 8 shows the empirical convergence results of SLRQA-NF-1, which indicate that even though the errors fluctuate at the early stage of iteration, they all converge in later iterations. We can observe that $\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1,1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1,2}\|_{\mathrm{F}}\rightarrow 0$. Hence, it follows from (87) that $|\|\widetilde{\mathbf{\dot{D}}}_{k+1}\||\rightarrow 0$. According to the update of $\mathbf{\dot{\Lambda}}_{k+1}$ and Lemma 4.16, it follows that $({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}}_{1},\mathbf{\dot{\Lambda}}_{2})$ converges to a stationary point of $\mathcal{L}_{\beta_{1},\beta_{2}}$ and hence a KKT point of the original problem in (89).
@@
Image Algorithm NCALR [25] QWSNM [48] QWNNM  [20] QSLRI [26] LRQA-1  [12] LRQA-2 [12] SLRQA-NF-1 SLRQA-NF-2 SLRQA-NF-3
$\chi=0.5$ Image1 26.48/0.85 28.16/0.86 25.25/0.60 31.16/0.95 30.87/0.89 30.25/0.87 32.32/0.94 31.26/0.94 32.48/0.94
Image2 25.55/0.90 24.56/0.87 23.55/0.81 30.89/0.85 26.63/0.91 26.20/0.90 28.04/0.94 27.54/0.94 28.17/0.94
Image3 26.75/0.91 27.87/0.90 25.23/0.75 29.65/0.97 28.37/0.90 27.94/0.89 30.40/0.96 29.65/0.97 30.61/0.97
Image4 28.47/0.93 28.56/0.92 31.16/0.85 36.39/0.99 35.99/0.97 35.51/0.97 37.69/0.99 36.27/0.98 37.91/0.99
Image5 33.46/0.94 34.23/0.91 30.09/0.87 33.73/0.98 33.27/0.96 32.72/0.95 34.82/0.98 33.75/0.98 34.98/0.98
Image6 30.26/0.90 31.03/0.87 22.01/0.61 26.56/0.92 25.27/0.79 25.72/0.80 28.25/0.93 26.62/0.91 27.97/0.92
Image7 27.48/0.91 28.65/0.92 24.96/0.75 28.84/0.94 27.59/0.87 27.86/0.88 27.82/0.93 28.92/0.94 29.33/0.94
Image8 28.12/0.93 29.42/0.94 25.17/0.86 30.03/0.97 29.62/0.95 29.06/0.94 31.06/0.97 30.16/0.97 31.22/0.97
Image9 29.83/0.94 28.89/0.95 26.02/0.94 30.99/0.98 30.12/0.98 29.56/0.97 31.58/0.98 31.09/0.98 31.71/0.98
Image10 29.47/0.95 28.43/0.93 26.18/0.94 30.69/0.98 29.86/0.98 29.20/0.97 31.28/0.98 30.84/0.98 31.42/0.98
$\chi=0.7$ Image1 27.39/0.84 27.41/0.85 23.05/0.52 27.29/0.89 25.97/0.74 25.47/0.72 28.40/0.88 27.30/0.88 28.62/0.89
Image2 25.66/0.87 24.11/0.85 20.13/0.67 24.19/0.89 22.51/0.79 22.44/0.79 25.05/0.90 24.30/0.89 25.22/0.90
Image3 26.44/0.93 24.12/0.92 21.50/0.63 26.09/0.93 23.69/0.77 23.67/0.77 26.87/0.93 26.00/0.92 27.10/0.94
Image4 31.32/0.96 30.47/0.95 28.34/0.80 31.89/0.97 31.07/0.93 30.44/0.92 33.37/0.97 31.80/0.97 33.63/0.98
Image5 26.55/0.91 27.46/0.92 25.75/0.77 29.58/0.95 28.53/0.88 27.85/0.86 30.95/0.96 29.68/0.95 31.14/0.96
Image6 23.77/0.74 20.57/0.75 18.13/0.44 22.78/0.82 20.28/0.57 19.86/0.55 23.78/0.81 22.70/0.80 24.00/0.83
Image7 24.01/0.83 25.43/0.85 22.40/0.67 26.24/0.90 24.62/0.78 24.58/0.78 26.63/0.88 26.19/0.89 26.57/0.88
Image8 26.08/0.93 25.45/0.92 22.34/0.79 26.34/0.94 24.87/0.87 24.54/0.87 27.41/0.94 26.42/0.94 27.60/0.95
Image9 22.36/0.87 24.15/0.93 23.33/0.90 27.83/0.97 26.13/0.94 25.69/0.94 28.56/0.97 27.81/0.97 28.74/0.97
Image10 23.14/0.91 23.76/0.92 23.22/0.90 27.54/0.96 25.95/0.95 25.34/0.94 28.46/0.97 27.58/0.96 28.64/0.97
$\chi=0.8$ Image1 19.43/0.79 23.16/0.81 20.77/0.43 25.05/0.85 22.87/0.61 22.76/0.60 26.08/0.84 24.54/0.79 26.30/0.85
Image2 17.65/0.80 22.47/0.85 17.72/0.54 22.34/0.84 19.44/0.65 20.27/0.69 23.36/0.86 22.25/0.84 23.53/0.86
Image3 22.65/0.84 23.45/0.86 18.71/0.53 23.98/0.90 20.52/0.65 21.35/0.69 24.91/0.90 23.61/0.88 25.13/0.91
Image4 28.41/0.93 27.48/0.92 25.48/0.72 29.09/0.95 27.75/0.87 27.20/0.85 30.74/0.96 28.73/0.94 31.04/0.96
Image5 26.08/0.88 26.45/0.89 22.52/0.65 27.10/0.93 25.53/0.80 25.05/0.78 28.70/0.93 27.09/0.92 28.97/0.94
Image6 18.04/0.65 19.76/0.74 15.41/0.30 20.82/0.74 17.93/0.45 14.65/0.31 21.61/0.74 20.10/0.70 21.82/0.76
Image7 22.18/0.80 23.16/0.79 20.67/0.60 24.79/0.87 22.81/0.71 22.64/0.71 25.60/0.85 24.53/0.86 25.42/0.86
Image8 21.38/0.84 24.74/0.90 19.88/0.71 24.26/0.91 21.69/0.79 21.93/0.80 25.31/0.92 23.91/0.90 25.50/0.92
Image9 24.07/0.89 25.19/0.91 20.78/0.84 25.89/0.95 23.43/0.91 23.24/0.91 26.74/0.96 25.56/0.95 26.97/0.96
Image10 22.92/0.90 23.79/0.92 20.48/0.84 25.52/0.95 23.10/0.90 22.85/0.90 27.67/0.96 25.33/0.95 26.91/0.96
Figure 8: The error convergence of SLRQA-NF for image inpainting problems.
@@ 6 Conclusion
In this paper, we propose a novel SLRQA for the color image process problems. Different from most existing models which only consider one or two properties from low-rankness, sparsity, and quaternion representation, SLRQA uses these properties all. Furthermore, SLRQA does not need an initial rank estimate. A PL-ADMM algorithm is proposed to solve the SLRQA and its global convergence is guaranteed. When the observation is noise-free, an SLRQA-NF of the limiting case of SLRQA is proposed. Subsequently, a PL-ADMM-NF algorithm is also proposed to solve the SLRQA-NF. Under a newly proposed assumption, the global convergence of PL-ADMM-NF is established. To the best of our knowledge, this is the first ADMM-type algorithm without the “range assumption” and still guarantees global convergence. Extensive experiments for color image denoising and inpainting problems have demonstrated the robustness and effectiveness of SLRQA and SLRQA-NF.
Data availability statement
The data that support the findings of this study are openly available at the following URL/DOI: https://github.com/dengzhanwang/SLRQA/tree/main.